14.6




Suppose a monopoly can produce any level of output it wishes at a constant marginal (and average) cost of $5 per unit. Assume the monopoly sells its goods in two different markets separated by some distance. The demand curve in the first market is given by




Q₁ = 55-P₁,




and the demand curve in the second market is given by




Q₂=70-2P₂.




a. If the monopolist can maintain the separation between the two markets, what level of output should be produced in each market, and what price will prevail in each market? What are total profits in this situation?




b. How would your answer change if it costs demanders only $5 to transport goods between the two markets? What would be




the monopolist's new profit level in this situation?

1
Expert's answer
2021-11-11T18:20:13-0500

a. If Q₁ = 55 - P₁, then P1 = 55 - Q1,

MR1 = 55 - 2Q.

MR = MC,

55 - 2Q1 = 5,

Q1 = 25 units,

P1 = 55 - 25 = 30.

If Q₂ = 70 - 2P₂, then P2 = 35 - 0.5Q2,

MR2 = 35 - Q2.

35 - Q2 = 5,

Q2 = 30 units.

P2 = 35 - 0.5×30 = 20.

"TP = (30 - 5)\u00d725 + (20 - 5)\u00d730 = 1,075."

b. If it costs demanders only $5 to transport goods between the two markets, then the good 2 only will be consumed.

c. The monopolist's new profit level in this situation will decrease.